Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - playing cards

Expert replies
by paridhi » Thu Jul 08, 2010 10:56 pm
Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

(A) 8/33

(B) 62/165

(C) 17/33

(D) 103/165

(E) 25/33

Can someone please tell me the approach to do this question?
Join the discussion
Source: — Problem Solving |

Hi,

Interesting problem ... Perhaps amongst the tough ones in GMAT.

Atleast 1 card repeating = total - No cards repeating.

Total card possibilities = 12 C 4 = 495

No card repeatings = 5 ways of achieving this

4 cards from deck1 (or)
3 cards from d1 and 1 of the other cards from d2 (or)
2 cards from d1 and 2 of the other cards from d2 (or)
1 card from d1 and 3 of the other cards from d2 (or)
4 cards from d2

=
6C4 +

6C3 times 3C1 +

6C2 times 4C2 +

6C1 times 5C3 +

6C4

= 240 ways

Atleast 1 card repeating = total - No cards repeating.

Atleast 1 card repeating = 495 - 240

P(the above event) = (495 - 240) / 495. Choice C

Look forward to a better method or pls highlight if there are some issues here.

Thanks.
Last edited by 4GMAT_Mumbai on Thu Jul 08, 2010 11:34 pm, edited 1 time in total.
Naveenan Ramachandran
4GMAT, Dadar(W) & Ghatkopar(W), Mumbai
Join the discussion

by Rahul@gurome » Thu Jul 08, 2010 11:26 pm
Let p be the probability that Bill finds no pair of cards that have the same value. Then probability that Bill finds at least one pair of cards that have the same value = 1- p
Total no. of ways in which 4 cards can be drawn from 12 cards = 12C4 = 495

Number of ways in which Bill finds no pair of cards with the same value:
There are 6 cards within a suit and we want to select 4 cards. So, this can be done in 6C4 = 15 ways.
For 15 values each card can be of either suit.
1st card can be chosen from 2 suits, similarly 2nd, 3rd & 4th card can be chosen from 2 suits. We can do this 2^4 = 16 ways
So, number of ways to choose 4 cards so that no pairs exist = 15×16 = 240

So, p = 240/495 = 16/ 33
Hence 1- p = 1-16/33 =17/33

[spoiler]The correct answer is (C).[/spoiler]
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by meshtrap » Thu Jul 08, 2010 11:33 pm
paridhi wrote:Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

(A) 8/33

(B) 62/165

(C) 17/33

(D) 103/165

(E) 25/33

Can someone please tell me the approach to do this question?
Since the problem asks us to find atleast one pair of cards that have the same value, a simpler approach would be to find the probability if none of the cards have the same value and subtract that number from 1.
P(A) = 1 - P(~A)

Thus, P(finding one pair) = 1 - P(not finding a single pair)

P(~Single Pair):

Calculate the denominator as follows:

4 cards can be chosen out of 12 cards in 12C4 = (12*11*10*9)/4! ways

Calculate the numerator as follows:

Since nCr = (nPr)/r!, calculate an ordered pair(permutation) and divide it by r!

1st card can be chosen in 12 ways
2nd------------------------ in 10 ways(if 2 is chosen, the other 2 in the deck cannot be chosen)
3rd------------------------- in 8 ways
4th--------------------------in 6 ways

Thus, numerator = (12*10*8*6)/4!

Thus, P(~Single pair) = num/den = (12*10*8*6)/(12*11*10*9) = 16/33

Thus P(Atleast one pair) = 1-16/33 = 17/33

Hope it helps,
meshtrap
Join the discussion

by paridhi » Thu Jul 08, 2010 11:40 pm
Thanks everyone for such good explanations.

OA Is C
Join the discussion